# Enhanced Muonization by Active-Sterile Neutrino Mixing in Protoneutron Stars

### v1

@@ V Discussion and conclusions
We have presented the effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\sim 10^{2}$ keV2 in the PNS. In addition to the obvious feedback on the $\nu_{\mu}$ lepton number, we have included for the first time, the feedback on the composition of the PNS. For our adopted mixing parameters, which are consistent with current constraints, we find that sterile neutrinos are dominantly produced through the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$. This production and the subsequent escape of $\bar{\nu}_{s}$ increase the $\nu_{\mu}$ lepton number, which in turn enhances muonization mainly through $\nu_{\mu}+n\to p+\mu^{-}$ and changes the number fractions $Y_{e}$, $Y_{\nu_{e}}$, $Y_{p}$, and $Y_{n}$ through chemical equilibrium. We find that diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ enhances the evolution of the number fractions of all the particles. In particular, such diffusion further enhances muonization.
In our adopted CCSN model 2017PhRvL.119x2702B , $Y_{\mu}$ reaches the peak value of 0.0566 at $r=8.97$ km at $t_{\rm pb}=1$ s with standard physics. By including $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\approx m_{s}^{2}\sim 10^{2}$ keV2, we have shown that enhanced muonization occurs in that model at larger radii beyond the original peak of $Y_{\mu}$ (Figs. 3 and 4). This enhancement is driven by the MSW resonances of $\bar{\nu}_{\mu}$ with $E_{R}\sim 4$–10 (8–20) MeV for $m_{s}=7.1$ (10) keV (Fig. 5), and is amplified by diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ (right panels of Figs. 3 and 4). While our treatment of such diffusion is approximate (e.g., with the same diffusion coefficient for all active neutrinos), it is adequate to demonstrate the above amplification effect and provide a reasonable estimate of this effect. In summary, we regard enhanced muonization as a robust result from $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\sim 10^{2}$ keV2 in the PNS.
@@
We have chosen to study the epoch of $t_{\rm pb}\approx 1$ s because the corresponding temperature and density profiles change so slowly that we can ignore such changes when studying the effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing for a duration of 0.1 s. On the other hand, the PNS conditions do not differ greatly for $t_{\rm pb}\sim 0.4$–2 s and the conditions at $t_{\rm pb}\sim 0.1$ s are also not qualitatively different, which suggests that our results can be extended to earlier and later times. With the help of muonization, neutrino-driven explosion occurs at $t_{\rm pb}\sim 0.24$ s in our adopted CCSN model 2017PhRvL.119x2702B . We expect that when $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\sim 10^{2}$ keV2 are included in CCSN simulations, enhanced muonization similar to that found here would occur at the explosion epoch, potentially making the explosion easier. However, due to the intrinsically dynamic nature of the explosion, a definitive conclusion can only be reached by implementing such mixing in CCSN simulations.
@@ ACKNOWLEDGMENT
We sincerely thank Daniel Kresse and Thomas Janka for providing the CCSN model used in this work. This work is supported in part by the National Science Foundation (Grant No. PHY-2020275), the Heising-Simons Foundation (Grant 2017-228), and the U.S. Department of Energy (Grant No. DE-FG02-87ER40328). *
@@ Appendix A Charged-current weak reaction rates
The evolution of $Y_{p}$ with time is governed by
$\dot{Y}_{p}$ $=\Gamma_{n\to p}Y_{n}-\Gamma_{p\to n}Y_{p}$
$=\Gamma_{n\to p}-(\Gamma_{n\to p}+\Gamma_{p\to n})Y_{p},$ (15)
where $\Gamma_{n\to p}$ and $\Gamma_{p\to n}$ are defined in terms of the rates for the charged-current weak reactions involving the electron and muon flavors [Eqs. (11a) and (11b)], and the second line follows from $Y_{n}+Y_{p}=1$. Regardless of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing, the rates for the charged-current weak reactions involving the electron flavor are very fast. Therefore, we can assume that $Y_{p}=\Gamma_{n\to p}/(\Gamma_{n\to p}+\Gamma_{p\to n})$ [Eq. (10d)] to very good approximation.
The rate for $\nu_{e}+n\to p+e^{-}$ is given by
@@
$\sigma_{\nu_{e}n}(E_{e^{-}})=\frac{G^{2}_{F}\cos^{2}\theta_{c}}{\pi}E_{e^{-}}p_{e^{-}}$ (17)
is the cross section, $\theta_{c}$ denotes the Cabibbo angle with $\cos^{2}\theta_{c}=0.97$,
$f_{\alpha}(E_{\alpha})=\frac{1}{\exp[(E_{\alpha}-\mu_{\alpha})/T]+1}\,,$ (18)
@@
The rate for $e^{+}+n\to p+\bar{\nu}_{e}$ is given by
$\Gamma_{e^{+}n}$ $=\frac{1}{n_{b}Y_{n}}\int_{0}^{\infty}\frac{p^{2}dp}{\pi^{2}}f_{n}(E_{n})[1-f_{p}(E_{p})]$
×$\times\int_{0}^{\infty}\frac{p_{e^{+}}^{2}dp_{e^{+}}}{\pi^{2}}f_{e^{+}}(E_{e^{+}})[1-f_{\bar{\nu}_{e}}(E_{\bar{\nu}_{e}})]\sigma_{e^{+}n}(E_{\bar{\nu}_{e}}),$ (24)
where
$\sigma_{e^{+}n}(E_{\bar{\nu}_{e}})=\frac{G^{2}_{F}\cos^{2}\theta_{c}}{2\pi}E_{\bar{\nu}_{e}}^{2}$ (25)
is the cross section, and $E_{\bar{\nu}_{e}}=E_{e^{+}}+m_{n}-m_{p}+U_{n}-U_{p}$.
The rate $\Gamma_{\mu^{+}n}$ for $\mu^{+}+n\to p+\bar{\nu}_{\mu}$ can be obtained by replacing quantities for $e^{+}$ ($\bar{\nu}_{e}$) with those for $\mu^{+}$ ($\bar{\nu}_{\mu}$) in the equations for $\Gamma_{e^{+}n}$.
The rate for $e^{-}+p\to n+\nu_{e}$ is given by
$\Gamma_{e^{-}p}$ $=\frac{1}{n_{b}Y_{p}}\int_{0}^{\infty}\frac{p^{2}dp}{\pi^{2}}f_{p}(E_{p})[1-f_{n}(E_{n})]$
×$\times\int_{E_{e^{-}}^{\rm th}}^{\infty}\frac{p_{e^{-}}^{2}dp_{e^{-}}}{\pi^{2}}f_{e^{-}}(E_{e^{-}})[1-f_{\nu_{e}}(E_{\nu_{e}})]\sigma_{e^{-}p}(E_{\nu_{e}}),$ (26)
where $\sigma_{e^{-}p}(E_{\nu_{e}})$ can be obtained by replacing $E_{\bar{\nu}_{e}}$ with $E_{\nu_{e}}$ in Eq. (25) for $\sigma_{e^{+}n}(E_{\bar{\nu}_{e}})$,
$E_{\nu_{e}}=E_{e^{-}}+m_{p}-m_{n}+U_{p}-U_{n},$ (27)
and $E_{e^{-}}^{\rm th}$ corresponds to $E_{\nu_{e}}=0$.
The rate $\Gamma_{\mu^{-}p}$ for $\mu^{-}+p\to n+\nu_{\mu}$ can be obtained by replacing quantities for $e^{-}$ ($\nu_{e}$) with those for $\mu^{-}$ ($\nu_{\mu}$) in the equations for $\Gamma_{e^{-}p}$.

### v2

@@ V Discussion and conclusions
We have presented the effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\sim 10^{2}$ keV2 in the PNS. In addition to the obvious feedback on the $\nu_{\mu}$ lepton number, we have included for the first time, the feedback on the composition of the PNS. For our adopted mixing parameters, which are consistent with current constraints, we find that sterile neutrinos are dominantly produced through the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$. This production and the subsequent escape of $\bar{\nu}_{s}$ increase the $\nu_{\mu}$ lepton number, which in turn enhances muonization mainly through $\nu_{\mu}+n\to p+\mu^{-}$ and changes the number fractions $Y_{e}$, $Y_{\nu_{e}}$, $Y_{p}$, and $Y_{n}$ through chemical equilibrium.
In our adopted CCSN model 2017PhRvL.119x2702B , $Y_{\mu}$ reaches the peak value of 0.0566 at $r=8.97$ km at $t_{\rm pb}=1$ s with standard physics. By including $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\approx m_{s}^{2}\sim 10^{2}$ keV2, we have shown that enhanced muonization occurs in that model at larger radii ($r\sim 12$–14.7 km) beyond the original peak of $Y_{\mu}$ (Figs. 3 and 4). This enhancement is driven by the MSW resonances of $\bar{\nu}_{\mu}$ with $E_{R}\sim 4$–10 (8–20) MeV for $m_{s}=7.1$ (10) keV (Fig. 7). We have ignored diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ in our calculations. Based on our adopted CCSN model with standard physics but with detailed treatment of such diffusion, the PNS is contracting between $t_{\rm pb}=1$ and 1.25 s and the resulting change in the density profile and more importantly, that in the temperature profile are most likely the dominant factors driving the evolution of composition profiles (see Fig. 2). However, changes of $Y_{\nu_{e}}$, $Y_{\nu_{\mu}}$, and especially $Y_{\mu}$ found in our calculations (assuming fixed temperature and density profiles over a duration of 0.1 s) are qualitatively different from those in the CCSN model with standard physics (see Figs. 3 and 4). Further, in the region of $r\sim 12$–14.7 km, the increase of $Y_{\mu}$ obtained from our calculations quantitatively exceeds the corresponding change in the standard CCSN model. Therefore, we regard enhanced muonization as a qualitatively robust result from $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\sim 10^{2}$ keV2 in the PNS, and expect that it still occurs when such mixing is incorporated in CCSN simulations. Such self-consistent simulations should be carried out to assess quantitatively the effects discussed here.
@@
We have chosen to study the epoch of $t_{\rm pb}\approx 1$ s because the corresponding temperature and density profiles change so slowly that we can ignore such changes when studying the effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing for a duration of 0.1 s. On the other hand, the PNS conditions do not differ greatly for $t_{\rm pb}\sim 0.4$–2 s and the conditions at $t_{\rm pb}\sim 0.1$ s are also not qualitatively different, which suggests that our results can be extended to earlier and later times. With the help of muonization, neutrino-driven explosion occurs at $t_{\rm pb}\sim 0.24$ s in our adopted CCSN model 2017PhRvL.119x2702B . We expect that when $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing with $\delta m^{2}\sim 10^{2}$ keV2 is included in CCSN simulations, enhanced muonization similar to that found here would occur at the explosion epoch, potentially making the explosion easier. However, due to the intrinsically dynamic nature of the explosion, a definitive conclusion can only be reached by implementing such mixing in CCSN simulations.
@@ ACKNOWLEDGMENT
We sincerely thank Daniel Kresse and Thomas Janka for providing the CCSN model used in this work. This work was supported in part by the National Science Foundation (Grant No. PHY-2020275), the Heising-Simons Foundation (Grant 2017-228), and the U.S. Department of Energy (Grant No. DE-FG02-87ER40328). *
@@ Appendix A Charged-current weak reaction rates
Under the so-called elastic approximation (the nucleons carry the same momentum), the rate for $\nu_{e}+n\to p+e^{-}$ is given by
@@
$\sigma_{\nu_{e}n}(E_{e^{-}})=\frac{G^{2}_{F}}{\pi}\cos^{2}\theta_{c}(f^{2}+3g^{2})E_{e^{-}}p_{e^{-}}$ (13)
is the cross section, $\theta_{c}$ denotes the Cabibbo angle with $\cos^{2}\theta_{c}\approx 0.95$, $f=1$ and $g\approx 1.27$ are weak coupling constants,
$f_{\alpha}(E_{\alpha})=\frac{1}{\exp[(E_{\alpha}-\mu_{\alpha})/T]+1}\,$ (14)
@@
By detailed balance, the rates for $e^{+}+n\to p+\bar{\nu}_{e}$, $\mu^{+}+n\to p+\bar{\nu}_{\mu}$, $e^{-}+p\to n+\nu_{e}$, and $\mu^{-}+p\to n+\nu_{\mu}$ are given by
$\Gamma_{e^{+}n}$ $=\frac{Y_{p}}{Y_{n}}\Gamma_{\bar{\nu}_{e}p}\exp\left(\frac{\mu_{\nu_{e}}+\mu_{n}-\mu_{p}-\mu_{e}}{T}\right),$ (20)
$\Gamma_{\mu^{+}n}$ $=\frac{Y_{p}}{Y_{n}}\Gamma_{\bar{\nu}_{\mu}p}\exp\left(\frac{\mu_{\nu_{\mu}}+\mu_{n}-\mu_{p}-\mu_{\mu}}{T}\right),$ (21)
$\Gamma_{e^{-}p}$ $=\frac{Y_{n}}{Y_{p}}\Gamma_{\nu_{e}n}\exp\left(\frac{\mu_{e}+\mu_{p}-\mu_{n}-\mu_{\nu_{e}}}{T}\right),$ (22)
$\Gamma_{\mu^{-}p}$ $=\frac{Y_{n}}{Y_{p}}\Gamma_{\nu_{\mu}n}\exp\left(\frac{\mu_{\mu}+\mu_{p}-\mu_{n}-\mu_{\nu_{\mu}}}{T}\right).$ (23)
The exponential factors in the above equations reduce to unity when the relevant particles are in chemical equilibrium. We have checked that the charged-current weak reaction rates are sufficiently fast, and therefore, chemical equilibrium is achieved to very good approximation.